OpenStudy (anonymous):

a manager wants to separate his 43 emploees into two teams. one team has two fewer memebers than half the other team. how many employees are in each team?

OpenStudy (texaschic101):

a = one team b = the other team a + b = 43 1/2b - 2 = a a + b = 43 (1/2b - 2) + b = 43 3/2b - 2 = 43 3/2b = 43 + 2 3/2b = 45 (3/2b)/(3/2) = 45/(3/2) b = 45/1 x 2/3 = 90/3 = 30 a + b = 43 a + 30 = 43 a = 43 - 30 a = 13 one team (a) has 13 employees and the other team (b) has 30 employees :)